A rectangular parking lot has a length that is 5 yards greater than the width. The area of the parking lot is 500 square yards. Find the length and the width.
Answer by ikleyn(52780)

.

x*(x+5) = 500


x^2 + 5x - 500 = 0

(x-20)*(x+25) = 0


x = 20 yards, the width.


ANSWER.  The width is 20 yards.  The length is 25 yards.

Solved.